Reports 1-1 of 1 Clear search Modify search
AOS (Beam Reducing Telescopes)
tomotada.akutsu - 22:06 Tuesday 11 December 2018 (7330) Print this report
Transmission lights at TMSX after locking of the Xarm cavity

Nakano, Akutsu

Checked the status of the both transmission light beams (green and IR), as the arm cavity became to lock stably.

  • The green spot positions were a few mm lowered. But I determined to leave the TMS-VIS as was; this could be still ok for the use, and if it will not be in the future, let's improve the alignment of TMS-VIS (Fig 1-5).
  • Then I detached the final iris on the TMS breadboard. During the work I accidentaly shaked the TMS-VIS (around 18:30?) and that can be confirmed by the LVDTs...
  • We could also see the focused IR spots with sensor cards! This could be the first stable spots that transmitted through the 3-km arm and shrinked by TMSX and we can seecheeky Fig 6,7, and 8. The IR spots seemed very circular, but please be noted that there is a window (100mm dia effective) in front of the BRT but the transmitted IR beam would have 35.3 mm radius. Anyway we will take a beam profiler tomorrow...

Due to the remove of the iris (?) all the DoFs outputs from the LVDTs changed its average numbers... especially in vertical and roll. The roll signals seems noisier than the otheres now, but I am not sure from when this noisier thing happened.

Images attached to this report
Search Help